A firm operating in a Perfectly Competitive Market has costs equal to C(Q) 7Q which of the following is the firms supply as function of price? b. Q 14P
Answer #1

A. P/14

Explanation: C(Q) = 7Q2

So, Marginal cost = d/dQ (7Q2) = 14Q.

In perfect competition, marginal cost = marginal revenie i.e. price

So, P = 14Q

i.e. Q = P/14